OpenStudy (anonymous):

-4n + 7 + 2n +1 Answer this question and get a medal!

OpenStudy (anonymous):

just give me the answer, i already know it

OpenStudy (the_fizicx99):

Combine like terms. -4n +2n 7+1 -2x + 8

OpenStudy (scorcher219396):

-2n+8

OpenStudy (the_fizicx99):

-2n*

OpenStudy (anonymous):

Ill give u a clue. It is eather 1, -3, 3 or 4

OpenStudy (the_fizicx99):

There's no = sign, but if you were to divide -2n /-2 and 8/-2 its -4..

OpenStudy (anonymous):

Oh, opps ment to put -4n + 7 + 2n =1 not +1 lol sorry typo :(

OpenStudy (anonymous):

hello? are u going to answer?

OpenStudy (the_fizicx99):

Yea, back.

OpenStudy (anonymous):

its a one digget answer

OpenStudy (the_fizicx99):

Same process, Combine terms. -4n + 2n 7 = 1 -7 -7 -2n = -6 /-2 /-2 n = 3

OpenStudy (anonymous):

Oh wow u get a medal! Correct!

OpenStudy (the_fizicx99):

This 6th grade math.. >.>

OpenStudy (amistre64):

we would prefer people not demand answers, not fall into the trap of giving an answer when one is demanded.

OpenStudy (the_fizicx99):

She said she knew the answer, though I should've left the last steps to her.. sorry.

OpenStudy (amistre64):

lol yeah, when someone says they already know the answer you should let them show their work to verify it :)

OpenStudy (the_fizicx99):

;~;